Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2007-11-02
ISS On-Orbit Status 11/02/07

Day 148 for Clayton Anderson. Flight Day 11 for STS-120/10A; Day 9 of Joint Ops.

ISS/Shuttle crew wake-up: 1:08am EDT. Sleeptime: 5:08pm (ISS), 5:38pm (Shuttle).

After crew wake-up, onboard activities moved briskly in continuation of preparations for tomorrow's EVA-4 (see Note below). (EMU (Extravehicular Mobility Unit) suit and EVA tool configuration were completed, as well as almost all hazard assessments. The SSRMS (Space Station Remote Manipulator System) grappled and handed off the OBSS (Orbiter Boom Sensor System) extension arm to the SRMS (Shuttle RMS) for overnight parking, after completing the necessary MT (Mobile Transporter) translations from WS8 (Worksite 8) to WS3 and back.

FE-2 Clay Anderson started his day with the IWIS (Internal Wireless Instrumentation System, first transferring the recorded data of the last IWIS structural dynamics data take (during P6 berthing) from the standard four RSUs (Remote Sensor Units) in Lab, Node, FGB & SM (Service Module) to the FS (File Server) laptop. The RSUs were then reset (programmed) for the next data take, later in the morning for Doug Wheelock's workout on the Shuttle ergometer at ~5:50am EDT. (The second Lab RSU with accelerometer for recording the Shuttle exercise is currently located near the Orbiter's cycle ergometer on the middeck.)

FE-1 Yuri Malenchenko had ~2.5 hrs scheduled for the periodic/long-term inspection of the pressure hull in the SM Working Compartment (RO), looking for any moisture, deposits, mold, corrosion and pitting behind panels 130, 134, 135, 138, 139 and underneath the TVIS treadmill.

Continuing the current round of the regular monthly preventive maintenance of RS (Russian Segment) ventilation systems, Yuri also spent time in the FGB cleaning interior panel ventilation grilles.

Additionally, Malenchenko conducted the regular task (currently daily) of checkout/verification of IP-1 airflow sensors in the various RS hatchways, including the SM- & FGB-to-Soyuz tunnels, the DC1-to-Progress and FGB-to-Node-1 passageway (this is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons).

CDR Whitson worked on the UOP-6 (Utility Outlet Panel 6) PS-120 Junction Box on reconfigurations for the LAN (Local Area Network) for Node-2 integration. (Tasks included adding a mass storage cable (W-cable) to PS(Power Strip)-120, switching HRF (Human Research Facility) laptop power from HRF-1 to the W-cable, and configuring the ER5 (EXPRESS Rack 5) laptop with a PS-28 connection from the HRF-PC.)

Continuing Node-2 outfitting for ~2.5 hrs, the CDR and Clay Anderson installed the cabling for the ISL (Integrated Station LAN) Router, to be strung and connected to the ISL later this month after Node-2 relocation to the Lab front end.

In Node-1/Unity, Peggy also inspected and cleaned the FDS (Fire Detection & Suppression) system's bacteria filter und SDs (smoke detectors). (There are a total of two SD units in Node-1 (NOD1D3-1 & NOD1D3-03). Inspection included the Kapton tape at the rear of vacuum for damage or wear. If repair was necessary, photo documentation was required.)

On the IATCS (Internal Active Thermal Control System in the Lab, the FE-2 disconnected the AmiA (Antimicrobial Applicator), relieved the pressure in its canister and stowed AmiA for return. Afterwards, Clayton removed low-flow jumpers from ITCS locations, performed a flush and took an OPA (Ortho-Phthalaldehyde) sample for return on the Shuttle. (AmiA introduced OPA, an antimicrobial agent, into the Lab IATCS coolant fluid. An NH 3 (ammonia) sample was not required.)

Yuri performed the routine servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. (Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers. Weekly SOZh reports (on Sundays) to TsUP/Moscow deal with number & dates of water and urine containers, counter readings of water consumption & urine collection, and total operating time of the POTOK air filtration system.)

Clay later worked on the IMS (Inventory Management System), updating/editing its standard "delta file", including locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

Malenchenko again had about 60 minutes for general crew orientation (station familiarization & acclimatization), as is standard daily rule for the first two weeks after starting station residency.

In preparation for tomorrow's EVA, the CDR set up two DCS-760 digital camera batteries for recharge in the Battery Charger in the Airlock (A/L).

At ~2:50am EDT, the FE-1 downlinked two PAO TV messages of greetings during VHF comm passes over RGS (Russian Ground Site) to TsUP/Moscow, one to a much-valued member of the Moscow Region Administration, the other to the Balashiha Military Prosecutor's Office on the 25 th Anniversary of its formation.

ISS crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (FE-1, FE-2/full-time), RED resistive exerciser (CDR, FE-1)and VELO bike with bungee cord load trainer (FE-1).

Afterwards, Peggy transferred the crewmembers' exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (Heart Rate Monitor) data, followed by their erasure on the HRM storage medium (done six times a week).

After completing A/L EL (Equipment Lock) configuration for the 4 th spacewalk tomorrow and conducting several extensive reviews and teleconferences of EVA-4 timeline, procedures, hazard avoidance, etc. with all crewmembers and ground specialists, the two spacewalkers, Scott Parazynski (EV1) and Doug Wheelock (EV2), will begin their "campout" in the "Quest" A/L, starting mask prebreathe at ~3:53pm, then closing hatches and initiating depressurization of the CL (Crewlock) from 14.7 to 10.2 psi. Sleep for the ISS crew begins at 5:08pm, for the Shuttle crew at ~5:38pm.

EVA-4 Overview: Primary focus of Scott & Wheels' outside venture will be inspection and repair of the torn P6 Solar Array Wing (SAW). Afterwards, if certain criteria are satisfied (see below), EV1 & EV2 will get out of the way and monitor the deployment of the wing. Subsequent get-ahead tasks include retrieval of two APFRs (Articulating Portable Foot Restraints, #3 & #7) for inspection, if time available and room in Airlock, mating of a PDGF (Power & Data Grapple Fixture) "horseshoe" connector at Node-2, and installation of the Node 2 HR (hand rail) #371 that was deferred from EVA-2. After the regular overnight Campout by the spacewalkers for denitrogenation/pre-breathe, egress is expected around 6:35am EDT tomorrow morning. EVA duration will again be about 6.5 hrs. (Repair Concept: There are two damaged areas on the SAW 4B inboard blanket: a large rip from the hinge between array segments ("bays") 35 & 36, and a small gap between bays 37 & 38. To stabilize the blanket in those areas, the affected bays will be "stitched" across the respective hinges (35/36 & 37/38) with five jerry-rigged wire clamps similar to dress shirt cufflinks, utilizing existing pin holes for inserting the cufflink anchors and requiring one-side access only. Because of the location of the pin holes, each of the "stitches" fortunately will run along spine-like rigidizing lines, called "stiffeners". Criteria for going ahead with the subsequent deployment of the 4B Array Mast include:

Snag on 4B right blanket eliminated; Array electrically safed by hinge wire being cut & end secured; Five stabilizers installed across hinge (priority from inboard to outboard); Guide wire cleared from creating additional hazards (grommet snag, FOD, free float); and Low tension mode during deployment visually confirmed and blanket stability monitored after each one-half bay deploy (i.e., stabilizers remaining in place, blanket deploying symmetrically, no additional snags, etc.).

Why 4B Panel Repair is required:

Array is believed to be stable for the near term;
However, it needs to be able to take loads over the long term;
Right now, about 3 feet of the 15-ft long hinge are missing, which is designed to distribute a load of ~70 lbs across it, i.e. there is a load concentration;
Therefore, the blanket could tear further over time;
If the blanket tears too much, the array cannot be stabilized and would have to be jettisoned (loss of ~20 kW of power);
Thus, P6 deployment is the number one priority for the rest of the mission.

No CEO (Crew Earth Observations) target uplinked for today.

ISS Orbit (as of this morning, 7:38am EDT (= epoch)):

Mean altitude -- 342.2 km

Apogee height -- 344.5 km

Perigee height -- 339.8 km

Period -- 91.38 min.

Inclination (to Equator) -- 51.64 deg

Eccentricity -- 0.0003515

Solar Beta Angle -- -62.8 deg (magnitude decreasing)

Orbits per 24-hr. day -- 15.76

Mean altitude loss in the last 24 hours -- 145 m

Revolutions since FGB/Zarya launch (Nov. 98) -- 51259

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use